OpenStudy (anonymous):

Solve the inequality. |r+3| > 7

OpenStudy (anonymous):

\[|r+3|\ge 7\]

OpenStudy (zzr0ck3r):

|a|>=b imples a>=b and a<= -b so solve r+3>=7 and r+3<= -7 r>=4 or r<= -10 so (-infinity,-10]U[4,infinity)